Does the following have a question have unique solution? Why?

[ 4x - 6y = -7
-2x + 3y =18

Answer by Alan3354(69443)

Does the following have a question have unique solution? Why?

[ 4x - 6y = -7
-2x + 3y =18
-----------------
No, it has no solution.
----
4x - 6y = -7
-2x + 3y =18 times -2 --> 4x - 6y = -36
-----------
It's inconsistent, says -7 = -36